OpenStudy (anonymous):

A canoeist pitches a 0.145 kg baseball at 21 m/s to a friend in another canoe. Both canoes are initially at rest, and the mass of each canoe including passenger is 120 kg. How fast do the canoes move apart after the second canoeist catches the ball?

OpenStudy (anonymous):

it is based on the conservation of momentum or we can also use newtons third law to every action there is an equal reaction m*v=M*V=>m,v are for base ball's mass and velocity and other for canon 0.145*21=120*V=>V=0.025375ms^-1 :)